# Unveiling Scoring Processes: Dissecting the Differences between LLMs and Human Graders in Automatic Scoring

### v1

@@ Abstract
Large language models (LLMs) have demonstrated strong potential in performing automatic scoring for constructed response assessments. While constructed responses graded by humans are usually based on given grading rubrics, the methods by which LLMs assign scores remain largely unclear. It is also uncertain how closely AI’s scoring process mirrors that of humans, or if it adheres to the same grading criteria. To address this gap, this paper uncovers the grading rubrics that LLMs used to score students’ written responses to science tasks and their alignment with human scores. We also examine whether enhancing the alignments can improve scoring accuracy. Specifically, we prompt LLMs to generate analytic rubrics that they use to assign scores and study the alignment gap with human grading rubrics. Based on a series of experiments with various configurations of LLM settings, we reveal a notable alignment gap between human and LLM graders. While LLMs can adapt quickly to scoring tasks, they often resort to shortcuts, bypassing deeper logical reasoning expected in human grading. We found that incorporating high-quality analytical rubrics designed to reflect human grading logic can mitigate this gap and enhance LLMs’ scoring accuracy. These results caution against the simplistic application of LLMs in science education and highlight the importance of aligning LLM outputs with human expectations to ensure efficient and accurate automatic scoring.
@@ 1 Introduction
In science education, educators develop scientific items to evaluate student abilities in logical thinking of understanding scientific phenomena [28, 19]. As automatic scoring helps the number of students to receive feedback on their responses in a timely fashion, it becomes a critical step in today’s science education [26]. Inspired by the huge progress in the field of natural language process, some researchers [22, 13, 27] propose to develop automatic scoring systems based on advanced large language models (LLMs). These LLMs have demonstrated a strong ability in quickly adapting to new downstream tasks (i.e., automatic scoring in this paper) with limited or even without any training samples [3]. This advancement significantly saves the cost of developing automatic scoring models and thus helps more students benefit from high-quality science education.
However, it is still unclear whether these LLM graders behave as human graders, raising the potential risks of applying them to border cases. To fill this gap, this study focuses on two research questions: (1) Is there an alignment gap between the human and LLM graders during the scoring reasoning steps? (2) If so, could filling this gap increase LLM graders’ scoring accuracy? While there are not many studies that answer these two questions, some attempt to improve the scoring performance of LLM graders. These works first find that LLMs cannot provide accurate scoring feedback as humans. In addition, some prompt engineering strategies, such as Chain-of-Thoughts [18, 10] and In-context Examples [3], improve the scoring accuracy of LLM-based systems. It is crucial to recognize that these studies find the gap by simply comparing the LLM-generated and human-provided categorical grades, without examining the real inherent process of “thoughts” of LLM graders during automatic scoring. That is, even if LLMs can provide grading levels exactly matching with human graders, it doesn’t necessarily mean the gap does not exist because LLMs may draw the same conclusion on a path different from the expectation of humans. These unexpected paths usually are referred to shortcuts [5, 17], indicating that LLMs find a superficial way to make predictions without following a reasonable logic chain as human experts.
@@
Our experiments span 12 different assessment items in science education, majorly covering topics in physics. Initial findings reveal notable differences between the LLM-generated rubrics and those crafted by human graders, indicating a significant alignment gap. Specifically, we observe that providing graded responses as examples cannot fill this gap, while introducing the holistic rubrics could help. Surprisingly, our further qualitative analysis suggests that providing LLMs with graded student responses does not inherently teach them to understand the question better; rather, it encourages LLM graders to use shortcuts in scoring. This observation against traditional recognition in automatic scoring that providing in-context graded responses could help LLM graders make more accurate grades. These findings caution against overly simplistic applications of LLMs in automatic scoring. Finally, our experiment demonstrates that integrating high-quality analytic rubrics can improve the performance of LLM graders in automatic scoring, highlighting the importance of addressing the misalignment gap to enhance their effectiveness.
@@ 2 Related Works
Research on leveraging LLMs for automatic scoring growth rapidly as this new perspective could provide consistent, less-bias, and personalized grading feedback for students compared with traditional grading methods [25, 8, 7, 11]. Early work focuses on fine-tuning LLMs for automatic scoring [12, 15], which typically provides accurate grading labels as human graders. Specifically, they may collect thousands of labeled student responses or a large corpus related to a specific assessment item for model training. While this procedure helps LLMs to adapt to the assessment items and provide high-quality feedback, it also raises a significant challenge in preparing such training data, which is costly and time-consuming. Other researchers [22] switch to explore whether it is possible to transform the automatic scoring task into one of the LLMs’ pre-training tasks. In this way, the development of an LLM-based automatic scoring system does not require training. Although this approach saves costs from both collecting datasets and training models, it also remains two critical issues. One is that such systems cannot provide highly accurate scoring labels since the LLMs have no chance to update their parameters to adapt to the automatic scoring task. Also, it leads to an ethical issue that whether those un-trained LLMs are reliable to serve our interested scoring tasks? Recently, researchers [13, 2, 20, 4, 24] have shown that the first issue could be partially solved by using more advanced LLMs or designing dedicate prompt engineering strategies. In particularlly, these studies indicate that providing more graded examples within the prompting templates or prompting LLMs to provide explanations along with their predicted scoring results could enhance LLMs to provide more accurate scoring results. However, their results reveal that those enhanced LLM graders could mimic human graders for those simpler assessment items, while they may fail to provide human-like feedback on those harder ones. In addition, the second issue about ethical concerns of this prompting-based approach is still unclear.
Since automatic scoring is a sensitive task and has been directly connected with the assessment and evaluation of un-adult students, it is crucial to know on what basis LLM performs the scoring tasks and how it concludes the final score. To achieve validity in the automatic scoring task, it becomes vital to have explainability in this automatic scoring procedure [23, 29]. To address this research gap, this study makes an attempt in providing the explainability to the automatic scoring task.
@@ 3.1 Datasets
This study experiments on a dataset by asking middle school students to describe scientific models accounting for science phenomena [28]. The dataset includes 7 assessment tasks, out of which two assessment tasks have binomial scoring rubrics, and the remaining five tasks have trinomial scoring rubrics. One example of these assessment tasks is: “Develop a model that includes a particle view of matter to predict how the average kinetic energy and the temperature of a substance changes when thermal energy is transferred to or from a sample.” Each assessment task could include multiple assessment items, and some ask the students to draw a figure to respond. By ignoring these items requiring visual responses, we finally selected 12 assessment items, each of which contains a question description (background and question), rubrics (analytical and holistic rubrics), and the number of scoring levels (binomial or trinomial). Here, holistic rubrics evaluate the student response as a whole and do not break the assessment down into different specific rules. It considers the overall student response and then grades the response based on the overall quality. However, analytic rubrics break down the assessment with multiple rules. Each of these rules of the assessment is evaluated separately. It provides detailed criteria based on which the assessment is graded. Each of the criteria has its own rating scale. In the end, all the criteria scores are then summed up to come to the total score or the grade of the student response. For each assessment task, we collected around 800 student responses [28]. Each student response is annotated according to the rubrics by humans into different levels, such as “Beginning”, “Developing”, or “Proficient”. We randomly select 100 labelled student responses from each assessment item with balanced grading levels, and use them to test the automatic scoring performance of LLM graders.
@@ 3.2 Experimental Designs
Table 1: Performance of analytic rubric generation and automatic scoring tasks.
Settings Analytic Rubric Generation Automatic Scoring
#Rules Pre. Rec. F1 Acc. (%)
No Analytic Rubric Guided (control) 0.00±0.00 0.000±0.00 0.000±0.00 0.000±0.00 34.83±11.50
Human Analytic Rubric Guided (comparison) 2.25±0.83 1.000±0.00 1.000±0.00 1.000±0.00 50.41±12.18
Analytic Rubric (one-shot) 3.25±1.09 0.525±0.33 0.701±0.34 0.580±0.32 49.17±13.28
Analytic Rubric (full-shot) 2.50±0.87 0.688±0.34 0.681±0.34 0.664±0.33 49.41±10.19
Analytic Rubric (full-shot) + Holistic Rubrics 2.75±1.16 0.782±0.31 0.778±0.28 0.752±0.28 54.58±${}_{\pm 9.01}\,\,$
Analytic Rubric (full-shot) + Graded Responses 3.33±1.43 0.326±0.33 0.528±0.36 0.350±0.28 48.41±10.17
  • #Rules: average number of rules for each analytic rubric. Pre.: average precision score of generated rules. Rec.: average recall score of human-written rules. Acc.: student’s responses scoring accuracy. We report the results over 12 assessments in the format of “avg.±${}_{\pm\text{std.}}$”.
The second aspect we studied was whether improving the alignment between LLMs and humans also benefits their scoring accuracy. To study this problem, we instruct LLMs to perform the traditional automatic scoring task by providing different analytic rubrics, including the human-written ones, and LLM-generated ones collected from the previous experiments. In our hypothesis, if the LLM is guided by a more accurate analytic rubric, its automatic scoring performance will also be better.
Our experiment used the Mixtral-8x7B-instruct [9] model as our subject LLM. Mixtral is a sparse mixture-of-expert language model, consisting of 47B parameters in total. As an open-sourced language model, it has surpassed those closed-source models under human evaluations [9], such as ChatGPT-3.5 Turbo [14], Claude-2.1 [21], Gemini Pro [6], and Llama-2 70B-Chat [1]. It was selected because of its superior performance and free accessibility. For reproduction in future work, we let Mixtral generate its responses by using the greedy search with a maximum 1024-tokens length constraint. To better instruct Mixtral for our downstream tasks, we applied the role-play prompting strategy [30]. Specifically, each round of conversation with the Mixtral started with the same system prompt, i.e., “The assistant is an impartial science educator working in a middle school. His job is working under the supervision of the User.” For the analytic rubric generation task, we provide a detailed instruction as “Your job is to provide the analytic rubric for a science item. The analytic rubric includes a minimum set of rules, each of which covers a specific required action, and their complete collection describes the requirements of the entire task.” Similarly, for the automatic scoring task, Mixtral is instructed with “Your job is to evaluate the quality of student responses strictly following the Analytic Rubric provided previously.
@@ 3.3 Analysis Methods
The experimental results were analyzed in quantitative sense. Specifically, to analyze the alignment gap between LLMs and humans, we measured the differences between the human-crafted and LLM-generated analytic rubrics with precision, recall, and F1 scores, where the precision score is the percentage of LLM-generated rules that match human-crafted ones, the recall score is the percentage of human-craft rules that have been mentioned by the LLM, and the F1 score is the geometric averaging of precision and recall scores. The F1 scores of certain settings are further evaluated by the Student T-Test with confidence level $\alpha=0.05$, where the null hypothesis is that the given two average scores are equal.
To analyze the benefits of filling the alignment gap, Spearman’s rank correlation coefficient between the analytic rubric F1 score and automatic scoring performance was referenced. If the Spearman’s correlation is significantly greater than zero, it means that if the LLM-generated scoring rubric resembles the human-crafted one, the higher accuracy of the LLM’s automatic scoring results compared to the ground truth - thus, we could say that filling the gap could help LLMs make more valid gradings.
Additionally, we conducted simple qualitative case studies on the failed analytic rules generated by LLMs to explore how different characteristics impact the alignment gap.
@@ 4 Experiments
4.1 Evaluating Alignment Gap via Analytic
Rubric Generation Task
@@ 4.1.1 Settings
Each generated analytic rubric includes multiple rules, which are separated by a special symbol “|||” for post-processing. Given a set of rules generated by Mixtral, we compared their meanings with all human written rules by using a fine-tuned RoBERTa-large [16]. This fine-tuned model returns a semantic similarity score between zero to one, where returning one indicates the given pair of texts share the exactly same meaning. We consider a generated rule to be correct if any human-written rule shares a 0.6 similarity with it, and then, we compute the precision score of LLM-generated rules by dividing the number of correct rules by the number of total generated rules. Similarly, a human-written rule is recalled by Mixtral if it shares at least 0.6 similarity with any of the LLM-generated rules, and then we compute the recall score of human-written rules by dividing the number of recalled rules by the number of total human-written rules. We finally compute the F1 score as $F1=(2\times p\times r)/(p+r)$, where $p$ and $r$ refer to the precision score and recall score, respectively. Ideally, if Mixtral understands the science items as like human educators, its generated rules match all human-written rules, resulting 1.0 precision, 1.0 recall, and 1.0 F1 scores. The average performances over 12 science items under different experimental settings are listed in Table 1.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated strong potential in performing automatic scoring for constructed response assessments. While constructed responses graded by humans are usually based on given grading rubrics, the methods by which LLMs assign scores remain largely unclear. It is also uncertain how closely AI’s scoring process mirrors that of humans or if it adheres to the same grading criteria. To address this gap, this paper uncovers the grading rubrics that LLMs used to score students’ written responses to science tasks and their alignment with human scores. We also examine whether enhancing the alignments can improve scoring accuracy. Specifically, we prompt LLMs to generate analytic rubrics that they use to assign scores and study the alignment gap with human grading rubrics. Based on a series of experiments with various configurations of LLM settings, we reveal a notable alignment gap between human and LLM graders. While LLMs can adapt quickly to scoring tasks, they often resort to shortcuts, bypassing deeper logical reasoning expected in human grading. We found that incorporating high-quality analytical rubrics designed to reflect human grading logic can mitigate this gap and enhance LLMs’ scoring accuracy. These results underscore the need for a nuanced approach when applying LLMs in science education and highlight the importance of aligning LLM outputs with human expectations to ensure efficient and accurate automatic scoring.
@@ 1 Introduction
In science education, educators develop scientific items to evaluate student abilities in logical thinking of understanding scientific phenomena [1, 2]. As automatic scoring helps students receive feedback to their responses in a timely fashion, it becomes a critical step in today’s science education [3, 4]. Inspired by the substantial progress in the field of natural language processing, some researchers [5, CoT, 6] propose to develop automatic scoring systems based on advanced large language models (LLMs). LLMs [7, 8, 9] are machine learning models trained to understand and generate human natural language. Modern LLMs are trained with a massive corpus from the Internet, and they can help humans with various daily tasks, such as writing emails, planning travel, and so on. Different from conventional machine learning models, these LLMs have demonstrated a strong ability to quickly adapt to new downstream tasks (i.e., automatic scoring in this paper) with limited or even without any training samples [10]. Therefore, these advanced LLMs are leveraged to assist human educators for various education purposes, such as tutoring [11, 12], assessment and evaluations [5, CoT], and peer-to-peer learning [6]. This advancement significantly saves the cost of developing automatic scoring models and thus helps more students benefit from high-quality science education.
However, it is still unclear whether LLM graders behave as human graders, raising the potential risks of applying them to border cases. Practically, if an LLM grader incorrectly assigns a poorer score to a student, it will hurt not only the student’s interest in studying but also the teacher’s confidence in using these systems. Also, if it incorrectly provides a higher score to a student, it could fail to provide prompt feedback to students and guide their further studying. To fill this gap, this study focuses on two research questions: (1) Is there an alignment gap between the human and LLM graders during the scoring reasoning steps? (2) If so, could filling this gap increase LLM graders’ scoring accuracy? While limited studies have answer these two questions, some attempt to improve the scoring performance of LLM graders. Although some prompt engineering strategies, such as Chain-of-Thoughts [13, 14] and In-context Examples [10], improve the scoring accuracy of LLM-based systems, most of these LLMs cannot provide as accurate scoring feedback as humans do. It is crucial to recognize that these studies find the gap by simply comparing the LLM-generated and human-provided categorical grades without examining the real inherent process of ”thoughts” of LLM graders during automatic scoring. That is, even if LLMs can provide grading that exactly match human graders, the gap may still exist because LLMs may draw the same conclusion on a path different from humans. These unexpected paths usually are referred to shortcuts [15, 16], indicating that LLMs find a superficial way to make predictions without following a reasonable logic chain as human experts.
@@
Figure 1: One exemplar assessment item from the datasets[17]. We present their human-crafted meta-information, including the task description, holistic rubrics, and analytic rubrics.
Task This task is measuring a student’s proficiency in the following: Develop a model that explains how particle motion changes when thermal energy is transferred to or from a substance without changing state. Shwan had 3 dishes of water at room temperature. She cooled one dish, causing thermal energy to transfer from that dish to surroundings. She kept the middle dish at room temperature. She transferred thermal energy into the third dish by heating it. Then Shwan dropped a red-coated chocolate candy into each dish. Construct a model that shows what is happening to water particles and red dye particles in each dish. Holistic Rubrics
- Proficient Level: Student develops a model that fully identifies both water and dye particles and their motions while describing that water molecules move faster at higher temperatures (and vice versa).
- Developing Level: Student develops a model that partially identifies both water and dye particles and their motion while describing that water molecules move faster at higher temperatures (and vice versa). - Beginning Level: Student does not at all develop a model that identifies both water and dye particles and their motion while describing that water molecules move faster at higher temperatures (Vice Versa). Analytic Rubrics
- Rule 1: Water and dye molecules move slowly when the water is cold, faster when at room temperature, and faster when the water is hot.
- Rule 2: The key identifies water and dye particles. - Rule 3: The key identifies the particle’s motion (faster/slower). - Rule 4: The answer is a meaningful sentence.
Our experiments span 12 different assessment items in science education, majorly covering topics in physics. Initial findings reveal notable differences between the LLM-generated rubrics and those crafted by human graders, indicating a significant alignment gap. Specifically, we observe that providing graded responses as examples cannot fill this gap while introducing holistic rubrics could help. Surprisingly, our further qualitative analysis suggests that providing LLMs with graded student responses does not inherently teach them to understand the question better; rather, it encourages LLM graders to use shortcuts in scoring. This observation against traditional recognition in automatic scoring that providing in-context graded responses could help LLM graders make more accurate grades. These findings caution against overly simplistic applications of LLMs in automatic scoring. Finally, our experiment demonstrates that integrating high-quality analytic rubrics can improve the performance of LLM graders in automatic scoring, highlighting the importance of addressing the misalignment gap to enhance their effectiveness.
@@ 2 Related Works
Research on leveraging LLMs for automatic scoring is growing rapidly as this new perspective could provide consistent, less-bias, and personalized grading feedback for students compared with traditional grading methods [18, 19, 20, 21]. Early work focuses on fine-tuning LLMs for automatic scoring [22, 23], which typically provides accurate grading labels as human graders. Specifically, they may collect thousands of labeled student responses or a large corpus related to a specific assessment item for model training[24]. While this procedure helps LLMs to adapt to the assessment items and provide high-quality feedback, it also raises a significant challenge in preparing such training data, which is costly and time-consuming. Other researchers [5] switch to explore whether it is possible to transform the automatic scoring task into one of the LLMs’ pre-training tasks. In this way, the development of an LLM-based automatic scoring system does not require training. Although this approach saves costs from both collecting datasets and training models, it also remains two critical issues. One is that such systems cannot provide highly accurate scoring labels since the LLMs have no chance to update their parameters to adapt to the automatic scoring task. Also, it leads to an ethical issue of whether those un-trained LLMs are reliable enough to serve our interested scoring tasks. Recently, researchers [CoT, 25, 26, 27, 28] have shown that the first issue could be partially solved by using more advanced LLMs or designing dedicated prompt-engineering strategies. In particular, these studies indicate that providing more graded examples within the prompting templates or prompting LLMs to provide explanations along with their predicted scoring results could enhance LLMs to provide more accurate scoring results. However, their results reveal that those enhanced LLM graders could mimic human graders for those simpler assessment items, while they may fail to provide human-like feedback on those harder ones. More importantly, the second issue about ethical concerns of this prompting-based approach is still unclear.
Since automatic scoring is a sensitive task directly connected with the assessment and evaluation of un-adult students, it is crucial to know on what basis LLM performs the scoring tasks and how it concludes the final score. To achieve validity in the automatic scoring task, it becomes vital to have explainability in this automatic scoring procedure [29]. Asking LLMs to provide a chain-of-thought explanation for their predictions is a practical way to explore their internal decision-making procedures [30, 31]. By leveraging this technique, researchers [CoT, 27] effectively collect the explanation while grading each student response, and the human educators evaluate the reliability of each individual feedback based on these explanations. However, this approach can not provide a full picture of the internal understanding of LLMs at an item level. To address this research gap, this study proposes to provide the explainability to the automatic scoring task by prompting LLMs to provide the rubrics of each assessment item. By further analyzing these LLM-generated rubrics, human educators can build a comprehensive recognition of the abilities of LLM graders over item levels.
@@ 3.1 Datasets
This study experiments on a dataset by asking middle school students to describe scientific models accounting for science phenomena [17, 32]. The dataset includes 7 assessment tasks, out of which two assessment tasks have binomial scoring rubrics, and the remaining five tasks have trinomial scoring rubrics. We provide one example of these assessment items in Figure 1. Each assessment task could include multiple assessment items, and some ask the students to draw a figure to respond. By ignoring these items requiring visual responses, we finally selected 12 assessment items, each of which contains a question description (background and question), rubrics (analytical and holistic rubrics), and the number of scoring levels (binomial or trinomial). Please note that human experts craft all the information (i.e., questions description, holistic and analytic rubrics, scoring levels) of the assessment items. Here, holistic rubrics evaluate the student response as a whole and do not break the assessment down into different specific rules. It considers the overall student response and then grades the response based on the overall quality. On the other hand, analytic rubrics break down the assessment with multiple rules, where each of these rules of the assessment is evaluated separately. It provides detailed criteria based on which the assessment is graded. Each of the criteria has its own rating scale. In the end, all the criteria scores are then summed up to come to the total score or the grade of the student response. For each assessment task, we collected around 800 student responses [32]. Each student response is annotated according to the rubrics by humans into different levels, such as “Beginning”, “Developing”, or “Proficient”. We randomly selected 100 labeled student responses from each assessment item with balanced grading levels and used them to test the automatic scoring performance of LLM graders.
@@ 3.2 Experimental Designs
[…]
Template-1: Generating analytic rubrics with oneshot examples. %Round-1: Instruction on generating analytic rubrics with role-play.
System: The agent is an impartial science educator working in a middle school. His job is working under the supervision of the User.
User: Your job is to provide the analytic rubric for a science item. Agent: I will begin to work on my job. %Round-2: Providing in-context examples for analytic rubric generation.
User:
[…]
%Round-3: Requesting analytic rubric from the target item.
User:
[…]
Template-2: Grading student responses with oneshot examples for an assessment item. %Round-1: Instruction on automatic scoring with role-play.
System: The agent is an impartial science educator working in a middle school. His job is working under the supervision of the User.
User: Your job is to evaluate the quality of the response provided by a student to a science item. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must classify the response on a scale of “Beginning”, “Developing,” and “Proficient” by strictly following this format: [[rating]], for example: “Rating: [[Beginning]]”. Refer to the CONTEXT and RUBRIC while rating. Agent: I will begin to work on my job. %Round-2: Providing in-context examples for automatic scoring.
User:
[…]
Ground truth score of the example student response.}}\}}
[…]
Template-1: Generating analytic rubrics with oneshot examples. %Round-3: Requesting automatic scoring for the testing student response.
User:
[…]
@@
The second aspect we studied was whether improving the alignment between LLMs and humans also benefits their scoring accuracy. To study this problem, we instruct LLMs to perform the traditional automatic scoring task by providing different analytic rubrics, including the human-written ones and LLM-generated ones collected from the previous experiments. In our hypothesis, if the LLM is guided by a more accurate analytic rubric (more aligned/matched with the human written ones), its automatic scoring performance will also be better. Here, since all assessment items are meticulously designed by human experts, we define these human-crafted analytic rubrics provided by the dataset as the ground-truth rubrics for our task. In addition, please note that a rubric leading to a better grading performance does not necessarily mean it is a correct one (please see further discussions on case studies in Section 4.2.2).
Our experiment used the Mixtral-8x7B-instruct [9] model as our subject LLM. Mixtral is a sparse mixture-of-expert language model consisting of 47B parameters in total. As an open-sourced language model, it has surpassed those closed-source models under human evaluations [9], such as ChatGPT-3.5 Turbo [33], Claude-2.1 [34], Gemini Pro [7], and Llama-2 70B-Chat [8]. It was selected because of its superior performance and free accessibility. In addition, we follow previous work [13, 35] to set the hyper-parameters $temperature=0.0$ and $top-p=0.01$ during the text generation process to ensure the reproducibility of our experiments. To better instruct Mixtral for our downstream tasks, we applied the role-play prompting strategy [36]. Specifically, each round of conversation with the Mixtral started with the same system prompt, i.e., “The assistant is an impartial science educator working in a middle school. His job is working under the supervision of the User.” For the analytic rubric generation task, we provide a detailed instruction as “Your job is to provide the analytic rubric for a science item. The analytic rubric includes a minimum set of rules, each of which covers a specific required action, and their complete collection describes the requirements of the entire task.” Similarly, for the automatic scoring task, Mixtral is instructed with “Your job is to evaluate the quality of student responses strictly following the Analytic Rubric provided previously.Figure 3 provide the basic templates for both tasks.
@@ 3.3 Analysis Methods
The experimental results were analyzed quantitatively. Specifically, to analyze the alignment gap between LLMs and humans, we measured the differences between the human-crafted and LLM-generated analytic rubrics with precision, recall, and F1 scores, where the precision score is the percentage of LLM-generated rules that match human-crafted ones, the recall score is the percentage of human-craft rules that have been mentioned by the LLM, and the F1 score is the geometric averaging of precision and recall scores. The F1 scores of certain settings are further evaluated by the Student T-Test with confidence level $\alpha=0.05$, where the null hypothesis is that the given two average scores are equal. Spearman’s rank correlation coefficient between the analytic rubric F1 score and automatic scoring performance was referenced to analyze the benefits of filling the alignment gap. If the Spearman’s correlation is significantly greater than zero, we could say there is a correlation between the human-LLM alignment and the automatic scoring performance. Additionally, we conducted simple qualitative case studies on the failed analytic rules generated by LLMs to explore how different characteristics impact the alignment gap.
@@ 4 Experiments
4.1 Evaluating the alignment gap via the analytic rubric generation task.
@@ 4.1.1 Settings
Each generated analytic rubric includes multiple rules, which are separated by a special symbol “||||||| | |” for post-processing. Given a set of rules generated by Mixtral, we compared their meanings with all human written rules by using a fine-tuned RoBERTa-large [37]. This fine-tuned model returns a semantic similarity score between zero and one, where returning one indicates the given pair of texts share the exact same meaning. We consider a generated rule to be correct if any human-written rule shares a 0.5 similarity with it, and then, we compute the precision score of LLM-generated rules by dividing the number of correct rules by the number of total generated rules. Similarly, Mixtral recalls a human-written rule if it shares at least 0.6 similarities with any of the LLM-generated rules, and then we compute the recall score of human-written rules by dividing the number of recalled rules by the number of total human-written rules. We finally compute the F1 score as $F1=2\times(p\times r)/(p+r)$, where $p$ and $r$ refer to the precision score and recall score, respectively. Ideally, if Mixtral understands the science items as human educators, its generated rules match all human-written rules, resulting in 1.0 precision, 1.0 recall, and 1.0 F1 scores. The average performances over 12 science items under different experimental settings are listed in Table 1.
Table 1: Performance of analytic rubric (A.R.) generation and grading tasks. The performance of the LLM-generated rubrics for the analytic rubric generation task (i.e., Precision, Recall, and F1) is evaluated against human-written rubrics.
Settings Analytic Rubric Generation Grading
#Rules Pre. Rec. F1 Acc. (%)
No A.R. (control) 0.00±0.00 0.000±0.00 0.000±0.00 0.000±0.00 34.83±11.50
Human A.R. (comparison) 2.25±0.83 1.000±0.00 1.000±0.00 1.000±0.00 50.41±12.18
One-shot A.R. 3.25±1.09 0.525±0.33 0.701±0.34 0.580±0.32 49.17±13.28
Full-shot A.R. 2.50±0.87 0.688±0.34 0.681±0.34 0.664±0.33 49.41±10.19
Full-shot A.R. + Holistic Rubrics 2.75±1.16 0.782±0.31 0.778±0.28 0.752±0.28 54.58±${}_{\pm 9.01}\,\,$
Full-shot A.R. + Graded Responses 3.33±1.43 0.326±0.33 0.528±0.36 0.350±0.28 48.41±10.17
  • A.R.: analytic rubric. One-shot: randomly selects one A.R. as an in-context example. Full-shot: uses the rest of A.R. as in-context examples. #Rules: average number of rules for each analytic rubric. Pre.: average precision score of generated rules. Rec.: average recall score of human-written rules. Acc.: scoring accuracy on student responses. We report the results in the format of “avg.±${}_{\pm\text{std.}}$”.